# Faithfulness-Aware Uncertainty Quantification for Fact-Checking the Output of Retrieval Augmented Generation

### v2

@@ Abstract
Large Language Models (LLMs) enhanced with external knowledge retrieval, an approach known as Retrieval-Augmented Generation (RAG), have shown strong performance in open-domain question answering. However, RAG systems remain susceptible to hallucinations: factually incorrect outputs that may arise either from inconsistencies in the model’s internal knowledge or incorrect use of the retrieved context. Existing approaches often conflate factuality with faithfulness to the retrieved context, misclassifying factually correct statements as hallucinations if they are not directly supported by the retrieval. In this paper, we introduce franq (Faithfulness-based Retrieval Augmented UNcertainty Quantification), a novel method for hallucination detection in RAG outputs. franq applies different Uncertainty Quantification (UQ) techniques to estimate factuality based on whether a statement is faithful to the retrieved context or not. To evaluate franq and other UQ techniques for RAG, we present a new long-form Question Answering (QA) dataset annotated for both factuality and faithfulness, combining automated labeling with manual validation of challenging examples. Extensive experiments on long- and short-form QA across multiple datasets and LLMs show that franq achieves more accurate detection of factual errors in RAG-generated responses compared to existing methods.
@@
Fact-Checking the Output of Retrieval Augmented Generation
[…]
@@ 1 Introduction
Large Language Models (LLMs) are increasingly being used for a wide range of tasks. However, LLMs are prone to generate plausible but factually incorrect generations, termed hallucinations, due to limited training data, ambiguity of input queries, architectural limitations, etc. Huang et al. (2025). External knowledge retrieval can partially help prevent factual errors Shuster et al. (2021). Retrieval Augmented Generation (RAG; Lewis et al. (2020)) which uses dynamically retrieved knowledge as input to generate LLM responses has been proposed as a way to address these limitations.
@@
However, despite the use of external knowledge, RAG still produces hallucinations Shi et al. (2023). Moreover, the use of retrieved information makes it challenging to detect hallucinations in LLM responses and determine their original source. The model becomes more confident in generating statements that appear in the retrieval, regardless of their correctness (Kim et al., 2024). The internal knowledge of the LLM and the retrieved information may contradict each other (Wang et al., 2024a, 2025). The retrieved information may even be erroneous, incomplete, or completely irrelevant to the query (Shi et al., 2023; Ding et al., 2024).
@@
Thus, an important question is how to define hallucination in a RAG system, given the interplay between the model’s internal knowledge and the retrieved context. One approach is to consider any content not directly supported by the retrieved context as a hallucination Niu et al. (2024). However, we argue that hallucination should be defined based on factual inaccuracies rather than strict contextual alignment. Specifically, a generated statement that originates from the LLM’s internal knowledge but lies outside the retrieved context should not be considered a hallucination if it is factually correct.
@@
Figure 1: franq illustration. Left: A user poses a question, and the RAG retrieves relevant documents and formulates an answer, potentially using information from the retrieved documents. Middle: The RAG output is decomposed into atomic claims. Right: The franq method assesses factuality by evaluating three components: faithfulness and factuality under both faithful and unfaithful conditions.
@@
To address this distinction, we differentiate between factuality and faithfulness. Faithfulness refers to whether the generated output is semantically entailed by the retrieved context, while factuality indicates whether the content is objectively correct Dziri et al. (2022); Yang et al. (2024). In the context of fact-checking of RAG output, detecting non-factual claims is more critical than identifying unfaithful ones, although assessing faithfulness can still indirectly aid in fact-checking.
@@
In this paper, we investigate the detection of non-factual statements produced by RAG using Uncertainty Quantification (UQ) techniques. We introduce franq (Faithfulness-based Retrieval Augmented UNcertainty Quantification), a novel method that first evaluates the faithfulness of the generated response and subsequently applies different UQ methods based on the outcome.
@@
We evaluate franq on both long-form and short-form QA tasks. For long-form QA, where answers include multiple claims, we assess factuality on the claim level. For that, we introduce a dataset with factuality annotations, combining automated labeling with manual validation of challenging examples. For short-form QA, where answers are brief, we treat each response as a single claim and evaluate it as a whole. We test our method on four QA datasets in this setting.
The key contributions of this work are:
@@
  • We develop a new UQ method for RAG, franq, that estimates uncertainty by first assessing faithfulness, and then enabling separate uncertainty quantification methods for faithful and unfaithful outputs; see Section 2.
@@
  • We develop a long-form QA factuality dataset for RAG. The dataset incorporates both factuality and faithfulness labels, and was build by combining automatic annotation with manual validation for difficult cases; see Section 3.
@@
  • We conduct comprehensive experimental validation on both long-form and short-form QA across several LLMs, demonstrating that franq significantly improves the detection of factual errors in RAG outputs compared to other approaches; see Section 4.
@@ 2 Uncertainty Quantification for RAG
Let $\mathbf{x}$ be the user query submitted to the RAG system. The system first retrieves $k$ passages, denoted by $\textbf{r}=\{r_{1},\dots,r_{k}\}$, from an external knowledge source using $\mathbf{x}$ as the query. The RAG model then employs a LLM to generate a greedy output $\mathbf{y}$, conditioned on both $\mathbf{x}$ and r.
Autoregressive LLMs produce text sequentially, generating one token at a time. At each step $t$, the model samples a token $y_{t}\sim p(\cdot\mid\mathbf{y}_{<t})$, where $\mathbf{y}_{<t}$ denotes the sequence of previously generated tokens. In the case of greedy decoding, this token is selected as the most likely outcome, i.e., $y_{t}=\arg\max_{y}{p(y\mid\mathbf{y}_{<t})}$. From $\mathbf{y}$, we extract a set of $l$ atomic claims, denoted as ${c_{1},\dots,c_{l}}$. Each claim $c_{i}$ is associated with a specific span of tokens, $\mathcal{S}(c_{i})$, which represents the indices of the tokens in $\mathbf{y}$ that correspond to this particular claim.
A claim $c$ is considered factually true if it is generally true, and false otherwise. A claim is deemed faithful with respect to the retrieved documents r if it is entailed by them, and unfaithful otherwise. While most current benchmarks for evaluating RAG outputs focus on evaluating faithfulness Dziri et al. (2022); Niu et al. (2024), our main objective is to assess the factuality of claims.
General baselines. A straightforward approach to hallucination detection is to concatenate $\mathbf{x}$ and r, and apply standard UQ techniques directly to this joint prompt. While this approach allows reuse of existing UQ methods, it disregards the structural distinction between the question and its retrieved context.
@@
Perplexity Fomicheva et al. (2020)
TokenSAR Duan et al. (2024)
Sample diversity Semantic Entropy Kuhn et al. (2023)
@@
Lexical Similarity Fomicheva et al. (2020)
Degree Matrix Lin et al. (2024)
Eccentricity Lin et al. (2024)
Sum of Eigenvalues Lin et al. (2024)
Number of Semantic Sets Lin et al. (2024)
Reflexive P(True) Kadavath et al. (2022)
Table 1: Summary of general baseline methods.
For example, a standard UQ baseline is the probability of a claim $c$ according to the LLM:
$p(c\mid\mathbf{x},\textbf{r})=\prod_{t\in\mathcal{S}(c)}p(y_{t}\mid\mathbf{x},\textbf{r},\mathbf{y}_{<t}).$ (1)
We select methods available in LM-Polygraph library Fadeeva et al. (2023) and summarize them in Table 1. More methods support short-form QA since the entire answer is assessed as a single claim, allowing the use of sampling-based techniques. For long-form QA, there are fewer applicable methods, as they must operate at the token level to aggregate information over claim spans.
2.1 Faithfulness-Based Retrieval Augmented uNcertainty Quantification (franq)
We introduce franq, a new approach for assessing the factuality of claims in RAG outputs by leveraging UQ techniques and explicitly treating $\mathbf{x}$ and r as separate inputs. The key idea is to first assess whether a generated claim is faithful to r and then apply different UQ strategies depending on the outcome. The overall factuality of a claim $c$ is estimated as:
@@ 2.1 Faithfulness-Based Retrieval Augmented uNcertainty Quantification (franq)
$P(c\text{ is unfaithful to {r}})\cdot P(c\text{ is true}\mid\text{unfaithful}).$
Here, we calculate $P(c\text{ is unfaithful to {r}})$ as $1-P(c\text{ is faithful to {r}})$. This decomposition integrates three probability components, each instantiated using a suitable UQ method:
@@
An overview of our approach is visually depicted in Figure 1. Detailed examples of franq applied to specific claims are given in Appendix E.
@@
We now describe the three components used in equation (2).
Faithfulness. To assess whether a claim $c_{i}$ is entailed by the retrieved evidence r, we use AlignScore, a semantic similarity metric based on RoBERTa fine-tuned for factual alignment Zha et al. (2023). AlignScore evaluates the consistency between a claim and its supporting context. A comparison with alternative methods is provided in Appendix C.1.
In long-form QA, we compare the claim $c_{i}$ with the concatenated retrieved documents r using AlignScore. For short-form QA, where the full answer $\mathbf{y}$ is treated as a single claim, we provide question context by calculating AlignScore on pairs of the form $(\mathbf{x}\circ\mathbf{y},r_{j})$, with $\circ$’ denoting string concatenation.
Factuality under unfaithful condition. When a claim $c$ is unfaithful (i.e., not entailed by r), it likely originates from the LLM’s internal knowledge. We estimate its factuality using the model’s own probability estimates. Specifically, we introduce the Parametric Knowledge UQ method, which computes the likelihood of $c$ based solely on the LLM’s parametric knowledge (Mallen et al., 2023) without providing retrieved evidence r:
@@
This approach is effective in the long-form QA setting (see Appendix C.2). However, for short-form QA, more alternative general UQ baselines are available, such as the ones employing sample diversity. We find that Sum of Eigenvalues Lin et al. (2024) better approximates factuality in this scenario (see Appendix C.2).
Therefore, we estimate the factuality of unfaithful claims using Parametric Knowledge in long-form QA and using Semantic Similarity in short-form QA.
Factuality under faithful condition. If a claim $c$ is evaluated as faithful, we can additionally assess whether it is entailed by the retrievals using a method different from AlignScore. To this end, we introduce the MaxNLI UQ method, which uses a pre-trained Natural Language Inference (NLI) model He et al. (2023) to assess whether a claim $c$ is entailed (‘e’) or contradicted (‘c’) by any of the retrieved passages. We report the maximum entailment confidence across all retrievals:
$\textrm{MaxNLI}(c)=\max_{j=\overline{1,k}}\frac{P\left(\text{NLI}(c,r_{j})=\text{`$e$'}\right)}{P\left(\text{NLI}(c,r_{j})\in\{\text{`$e$'},\text{`$c$'}\}\right)},$ (4)
where probabilities are taken from the NLI model’s predicted class distribution. As with AlignScore, for long-form QA, we apply the NLI model to each individual claim, and for short-form QA, the premise is formed by concatenating the question $\mathbf{x}$ with the answer $\mathbf{y}$.
We later find that MaxNLI performs well for long-form QA, making it a reasonable choice for estimating factuality under the faithful condition, however, the specific choice of method for this component has limited impact on overall franq performance (see Appendix D.1).
Again, for short-form QA, we find alternative baselines more suitable, particularly Semantic Entropy Kuhn et al. (2023), which better captures uncertainty in this scenario (see Appendix C.2).
Therefore, we estimate the factuality for faithful claims with MaxNLI for long-form QA, and Semantic Entropy for short-form QA.
@@
where we use AlignScore to estimate faithfulness probability $P_{\text{faithful}}$ and two UQ methods, $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$, selected based on empirical performance for long-form and short-form scenarios. For long-form QA, we use MaxNLI (4) for $\text{UQ}_{\text{faith}}$ and Parametric Knowledge (3) for $\text{UQ}_{\text{unfaith}}$. For short-form QA, we use Semantic Entropy Kuhn et al. (2023) for $\text{UQ}_{\text{faith}}$ and Sum of Eigenvalues Lin et al. (2024) for $\text{UQ}_{\text{unfaith}}$.

### v3

# Faithfulness-Aware Uncertainty Quantification for Fact-Checking the Output of Retrieval-Augmented Generation

@@ Abstract
Large Language Models (LLMs) enhanced with retrieval, an approach known as Retrieval-Augmented Generation (RAG), have achieved strong performance in open-domain question answering. However, RAG remains prone to hallucinations: factually incorrect outputs may arise from inaccuracies in the model’s internal knowledge and the retrieved context. Existing approaches to mitigating hallucinations often conflate factuality with faithfulness to the retrieved evidence, incorrectly labeling factually correct statements as hallucinations if they are not explicitly supported by the retrieval. In this paper, we introduce franq, a new method for hallucination detection in RAG outputs. franq applies distinct uncertainty quantification (UQ) techniques to estimate factuality, conditioning on whether a statement is faithful to the retrieved context. To evaluate franq and competing UQ methods, we construct a new long-form question answering dataset annotated for both factuality and faithfulness, combining automated labeling with manual validation of challenging cases. Extensive experiments across multiple datasets, tasks, and LLMs show that franq achieves more accurate detection of factual errors in RAG-generated responses compared to existing approaches.
@@
Fact-Checking the Output of Retrieval-Augmented Generation
Ekaterina Fadeeva1 $\diamondsuit$   Aleksandr Rubashevskii2 $\diamondsuit$   Dzianis Piatrashyn2 Roman Vashurin2 Shehzaad Dhuliawala1 Artem Shelmanov2 Timothy Baldwin2 Preslav Nakov2   Mrinmaya Sachan1   Maxim Panov2 1ETH Zürich  2MBZUAI {efadeeva,sdhuliawala,msachan}@ethz.ch {aleksandr.rubashevskii,preslav.nakov,maxim.panov}@mbzuai.ac.ae
@@ 1 Introduction
Large Language Models (LLMs) are increasingly employed across a wide range of tasks. However, LLMs are prone to generating plausible but factually incorrect generations, a phenomenon known as hallucination, arising from factors such as insufficient training data coverage, input ambiguity, and architectural constraints Huang et al. (2025). Retrieval-Augmented Generation (RAG; Lewis et al., 2020) mitigates this issue by incorporating dynamically retrieved external knowledge into the generation process, which can partially mitigate factual inaccuracies Shuster et al. (2021).
@@
However, RAG systems still produce hallucinations Shi et al. (2023). Moreover, the use of retrieved information makes it more challenging to detect hallucinations and to determine their original source. Models become more confident in generating statements that appear in the retrieval, regardless of their factual correctness (Kim et al., 2025). At the same time, the retrieved passages themselves may be erroneous, incomplete, or completely irrelevant with respect to the query (Shi et al., 2023; Ding et al., 2024). Conversely, even when retrieval is accurate, inconsistencies can emerge between the model’s internal knowledge and the retrieved data (Wang et al., 2024a, 2025).
@@
Thus, an important question is how to define hallucination in RAG, given the interplay between the model’s internal knowledge and the retrieved context. One approach is to consider any content that is not directly supported by the retrieved context as a hallucination Niu et al. (2024). However, we argue that hallucination should be defined based on factual inaccuracies rather than strict contextual alignment. Specifically, a generated statement that originates from the LLM’s internal knowledge but lies outside the retrieved context should not be considered a hallucination if it is factually correct.
@@
Figure 1: franq illustration. Left: A user poses a question, and the RAG retrieves relevant documents and formulates an answer, potentially using information from the retrieved documents. Middle: The RAG output is decomposed into atomic claims. Right: The franq method assesses factuality by evaluating three components: (1) faithfulness, (2) factuality under faithful condition, and (3) factuality under unfaithful condition.
@@
To address this distinction, we differentiate between factuality and faithfulness. Faithfulness refers to whether the generated output is semantically entailed by the retrieved context, while factuality indicates whether the content is objectively correct Maynez et al. (2020); Dziri et al. (2022); Yang et al. (2024). For RAG fact-checking, detecting non-factual claims is more critical than identifying unfaithful ones. This distinction disentangles two core RAG failure modes: (i) hallucinations caused by erroneous grounding in the retrieved context, and (ii) factual errors stemming from the model’s internal knowledge Zhou et al. (2024).
@@
In this paper, we investigate the detection of non-factual statements produced by RAG using Uncertainty Quantification (UQ) techniques. We introduce franq (Faithfulness-aware Retrieval Augmented UNcertainty Quantification), a novel method that first evaluates the faithfulness of the generated response and subsequently applies different UQ methods based on the outcome. With this separation, franq tailors its strategy to the specific RAG failure mode: whether it originates from retrieval grounding or from the model’s own knowledge.
@@
We evaluate franq on both long- and short-form question answering (QA) tasks. For long-form QA, where answers include multiple claims, we assess factuality at the claim level and introduce a new dataset with factuality annotations, combining automated labeling with manual validation. For short-form QA, we test our method on four QA datasets and treat each response as a single claim.
Our key contributions are as follows.
@@
  • We develop a new UQ method for RAG, franq, that estimates uncertainty by first assessing faithfulness, and then using uncertainty quantification methods for faithful and unfaithful outputs; see Section 2.
@@
  • We develop a long-form QA factuality dataset for RAG. The dataset incorporates both factuality and faithfulness labels, and was built by combining automatic annotation with manual validation for difficult cases; see Section 3.
@@
  • We conduct comprehensive experiments on both long- and short-form QA with several LLMs, demonstrating that franq improves the detection of factual errors in RAG outputs compared to other approaches; see Section 4.
@@ 2 Uncertainty Quantification for RAG
Let $\mathbf{x}$ be the user query submitted to the RAG system. The system retrieves $k$ passages denoted by $\textbf{r}=\{r_{1},\dots,r_{k}\}$, from an external knowledge source using $\mathbf{x}$ as the query. The RAG system then uses an LLM to generate an output $\mathbf{y}$, conditioned on both $\mathbf{x}$ and r.
Autoregressive LLMs produce text sequentially, generating one token at a time. At each step $t$, the model samples a token $y_{t}\sim p(\cdot\mid\mathbf{y}_{<t},\mathbf{x},\textbf{r})$, where $\mathbf{y}_{<t}$ denotes the sequence of previously generated tokens. In the case of greedy decoding, this token is selected as the most likely outcome, i.e., $y_{t}=\arg\max_{y}{p(y\mid\mathbf{y}_{<t},\mathbf{x},\textbf{r})}$. From $\mathbf{y}$, we extract a set of $l$ atomic claims denoted as $c_{1},\dots,c_{l}$}. Each claim $c_{i}$ is associated with a specific span of tokens, $\mathcal{S}(c_{i})$, which represents the indices of the tokens in $\mathbf{y}$ that correspond to this particular claim.
A claim $c$ is considered factually true if it is generally true, and false otherwise. A claim is deemed faithful with respect to the retrieved documents r, if it is entailed by them, and unfaithful otherwise. While most current benchmarks for evaluating RAG outputs focus on evaluating faithfulness Dziri et al. (2022); Niu et al. (2024), our main objective is to assess the factuality of claims.
General baselines. A straightforward approach to hallucination detection is to apply standard UQ methods to the LLM output conditioned on the joint prompt containing both the user query $\mathbf{x}$ and the retrieved context r. However, this strategy ignores the structural asymmetry between $\mathbf{x}$ and r.
@@
CCP Fadeeva et al. (2024)
Reflexive P(True) Kadavath et al. (2022)
Sample diversity Lexical Similarity Fomicheva et al. (2020)
Degree Matrix Lin et al. (2024)
Sum of Eigenvalues Lin et al. (2024)
Semantic Entropy Kuhn et al. (2023)
@@
Table 1: Summary of UQ methods used as baselines.
As an illustrative example, a common UQ baseline is to estimate the negative log-probability of a claim $c$ under the model distribution:
$U(c\mid\mathbf{x},\textbf{r})=-\sum_{t\in\mathcal{S}(c)}\log p(y_{t}\mid\mathbf{x},\textbf{r},\mathbf{y}_{<t}).$ (1)
Table 1 summarizes several other UQ methods that can be applied in this general baseline setting.
2.1 Faithfulness-aware Retrieval Augmented uNcertainty Quantification (franq)
We introduce franq, a new approach for assessing the factuality of claims in RAG outputs by leveraging UQ and explicitly treating $\mathbf{x}$ and r as separate inputs. The key idea is to first assess whether a generated claim is faithful to r and then apply different UQ methods depending on the outcome. This yields the following decomposition of the probability that a claim $c$ is true:
@@ 2.1 Faithfulness-aware Retrieval Augmented uNcertainty Quantification (franq)
$P(c\text{ is unfaithful to {r}})\cdot P(c\text{ is true}\mid\text{unfaithful}),$
where $P(c\text{ unfaithf. to {r}})=1-P(c\text{ faithful to {r}})$. This decomposition isolates three probability components, each of which we approximate using specialized techniques described in Section 2.2:
@@
An overview of franq is visually depicted in Figure 1, and illustrative examples applied to individual claims are provided in Appendix G.
@@
We now describe the three components in equation (2).
Faithfulness. To determine the degree to which a claim $c_{i}$ is entailed by the retrieved evidence r, we use AlignScore, a RoBERTa-based similarity metric fine-tuned for factual alignment Zha et al. (2023). AlignScore is specifically designed to measure factual consistency between a claim and context evidence, making it well suited for claim-level faithfulness estimation in RAG. Importantly, AlignScore yields well-calibrated continuous faithfulness estimates rather than near-binary decisions; in practice, many claims exhibit intermediate values due to partial or implicit grounding. We analyze the distribution, calibration, and alternative faithfulness estimators in Appendix C.
In long-form QA, we apply AlignScore to each claim–retrieval pair $(c_{i},\textbf{r})$ to get the faithfulness estimate for claim $c_{i}$. In short-form QA, where the entire answer $\mathbf{y}$ is treated as a single claim, we prepend the question context and instead evaluate AlignScore on $(\mathbf{x}\circ\mathbf{y},\textbf{r})$, with $\circ$’ denoting string concatenation.
Factuality under unfaithful condition. When a claim $c$ is unfaithful (not entailed by r), it originates from the LLM’s internal knowledge. In this case, we estimate factuality using the model’s probability estimates, avoiding distributional shifts arising from conditioning on retrieved context r. Specifically, we introduce a Parametric Knowledge method, which computes the likelihood of $c$ based solely on the LLM’s parametric knowledge (Mallen et al., 2023) without the retrieved evidence r:
@@
This method does not require generating new responses; instead, it reuses the original tokens and performs a forward pass through the LLM with the retrieved evidence removed from the input.
In long-form QA, we find that Parametric Knowledge provides an effective estimate of factuality for unfaithful claims (see Section 2). In contrast, short-form QA admits a wider range of general UQ baselines, including methods based on sample diversity (see Table 1). In this setting, we observe that the Sum of Eigenvalues method Lin et al. (2024) offers a better approximation of factuality (see Section 2). Therefore, we estimate the factuality of unfaithful claims using Parametric Knowledge in long-form QA and Sum of Eigenvalues in short-form QA.
Factuality under faithful condition. When a claim $c$ is assessed as faithful to r, the LLM may still fail to apply that evidence correctly to the user query. For example, the LLM may simply choose one of the entities mentioned in r, producing a faithful but incorrect answer to the query $\mathbf{x}$.
To account for such errors, in long-form QA, we estimate uncertainty within the faithful branch using a simple Max Claim Probability baseline, $p(c\mid\mathbf{x},\textbf{r})$. In short-form QA, alternative baselines are more suitable, particularly Semantic Entropy Kuhn et al. (2023), which better captures uncertainty in this scenario (see Section 2).
Therefore, we estimate the factuality for faithful claims with Max Claim Probability for long-form QA, and Semantic Entropy for short-form QA.
@@
where we use AlignScore to estimate faithfulness probability $P_{\text{faithful}}$ and two UQ methods, $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$, selected based on empirical performance for long- and short-form QA scenarios. For long-form QA, we use Max Claim Probability (1) for $\text{UQ}_{\text{faith}}$ and Parametric Knowledge (3) for $\text{UQ}_{\text{unfaith}}$. For short-form QA, we use Semantic Entropy Kuhn et al. (2023) for $\text{UQ}_{\text{faith}}$ and Sum of Eigenvalues Lin et al. (2024) for $\text{UQ}_{\text{unfaith}}$.
We consistently apply the same uncertainty methods across all datasets within each QA setting (short- and long-form), and select uncertainty techniques only based on the nature of the task (using token-level likelihoods for long-form QA and sampling-based metrics for short-form QA).